OpenStudy (anonymous):

f(x) = -2x-7 and g(x) = -4x + 3. Find (f o g) (-5)

OpenStudy (anonymous):

(f o g)(x)=f(g(x))=-2(-4x+3)-7=8x-6-7=8x-13 f(g(-5))=8(-5)-13=-53

OpenStudy (anonymous):

thank you !! @ASAAD123